OpenStudy (anonymous):

(-1,20) perpendicular to the graph of 4x+20y=15

OpenStudy (amistre64):

x +5y = 99 perhaps?

OpenStudy (amistre64):

nope.... -5x +y = 25

OpenStudy (anonymous):

y=mx+c 20y=-4x+15 divide all through by 20 gradient will be -1/5 perpendicular is normally if you get the product it becomes -1 thus : -1/5 * x=-1 so x=5 so amistres is correct